What is the required return on the preferred shares of TH Incorporated? The price of the...

  1. What is the required return on the preferred shares of TH Incorporated? The price of the TH preferred share is $75, the par value of the share is $85 and the dividend rate is 11%.
  1. 13.2%
  2. 9.7%
  3. 11%
  4. 12.5%
Answer #1

Annual dividend=85*11%=$9.35

Required rate of return=Annual dividend/Current value

=9.35/75

which is equal to

=12.5%(Approx)